A solution contains Pb(BrO3)2 and PbS. Propose a way in which you could separate Pb(BrO3)2 (Ksp: 7.9 × 10-6) from PbS (Ksp: 3 × 10-28) contained in the solution.
Answer #1

A solution contains Pb(BrO3)2 and PbS. Pb(BrO3)2 (Ksp: 7.9 × 10-6) and PbS (Ksp: 3 × 10-28) .

we have :

Ksp (Pb(BrO3)2) = [Pb2+] [BrO3-]2 =7.9 × 10-6    ===>   S = (7.9 × 10-6 )^1/3 = 0.020 M

Ksp (PbS) = [Pb2+] [S2-] = 3 × 10-28 ===>   S = (3 × 10-28 )^1/2 = 1.732× 10-14 M

If we add, ~ 0.001 M of [Pb2+] ions in form of soluble salt (Better, Pb(BrO3)2 ) we will have all PbS (s) in solid or precipitate form, and we will have Pb(BrO3)2 reamining in solution as soluble form.